if f(1)=3 and f(n)= -2f(n-1) +1, then f(5) = ???
Answer by jim_thompson5910(35256)

f(1)=3 is given
--------------------
f(n)= -2f(n-1) +1
f(2)= -2f(2-1) +1 ... replace n with 2
f(2)= -2f(1) +1
f(2)= -2*3 +1 ... replace f(1) with 3
f(2) = -5

--------------------
f(n)= -2f(n-1) +1
f(3)= -2f(3-1) +1 ... replace n with 3
f(3)= -2f(2) +1
f(3)= -2*(-5) +1 ... replace f(2) with -5
f(3) = 11

--------------------
f(n)= -2f(n-1) +1
f(4)= -2f(4-1) +1 ... replace n with 4
f(4)= -2f(3) +1
f(4)= -2*(11) +1 ... replace f(3) with 11
f(4) = -21


--------------------
f(n)= -2f(n-1) +1
f(5)= -2f(5-1) +1 ... replace n with 5
f(5)= -2f(4) +1
f(5)= -2*(-21) +1 ... replace f(4) with -21
f(5) = 43
-------------------------------------------------------

Summary:

f(1) = 3
f(2) = -5
f(3) = 11
f(4) = -20
f(5) = 43

Notice how we had to build our way up to get to f(5). We started with f(1) which was given. Then we used that to find f(2). Then we used f(2) to find f(3), etc etc until got to f(5). This is a recursive function. Each term is found by using the previous term (except for the given f(1) term). It isn't so bad for small values of n, but gets somewhat tedious to do by hand for large values of n.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------